### v4

@@ 2.2 Computational Complexity
We assume that the reader is familiar with standard complexity classes such as $\mathsf{L}$, $\mathsf{NL}$, $\mathsf{NP}$, and $\mathsf{EXP}$. For a standard reference on circuit complexity, see [Vol99]. We consider Boolean circuits using the AND, OR, NOT, and Majority, where $\textsf{Majority}(x_{1},\ldots,x_{n})=1$ if and only if $\geq n/2$ of the inputs are $1$. Otherwise, $\textsf{Majority}(x_{1},\ldots,x_{n})=0$. In this paper, we will consider DLOGTIME-uniform circuit families $(C_{n})_{n\in\mathbb{N}}$. For this, one encodes the gates of each circuit $C_{n}$ by bit strings of length $O(\log n)$. Then the circuit family $(C_{n})_{n\geq 0}$ is called $\mathsf{DLOGTIME}$-uniform if (i) there exists a deterministic Turing machine that computes for a given gate $u\in\{0,1\}^{*}$ of $C_{n}$ ($|u|\in O(\log n)$) in time $O(\log n)$ the type of gate $u$, where the types are $x_{1},\ldots,x_{n}$, NOT, AND, OR, Majority, or oracle gates, and (ii) there exists a deterministic Turing machine that decides for two given gates $u,v\in\{0,1\}^{*}$ of $C_{n}$ ($|u|,|v|\in O(\log n)$) and a binary encoded integer $i$ with $O(\log n)$ many bits in time $O(\log n)$ whether $u$ is the $i$-th input gate for $v$.
@@
[…]
  • The AND and OR gates take exactly $2$ inputs. That is, they have fan-in $2$.
@@
[…]
Furthermore, for a language $L$ the class $\mathsf{AC}^{k}(L)$, apart from Boolean gates, also allows oracle gates for $L$ (an oracle gate outputs $1$ if and only if its input is in $L$). If $K\in\mathsf{AC}^{k}(L)$, then $K$ is said to be $\mathsf{AC}^{k}$ Turing reducible to $L$. Finally, for some complexity class $\mathcal{C}$ denote $\mathsf{AC}^{k}(\mathcal{C})$ to be the set of decision problems that are $\mathsf{AC}^{k}$-Turing reducible to problems in $\mathcal{C}$. Be aware that here we follow the notation of [Vol99], which is different from [Wag10, GL23] (where $\mathsf{AC}^{k}(\mathcal{C})$ is used to denote composition of functions).
@@
$\textsf{NC}^{k}\subseteq\mathsf{SAC}^{k}\subseteq\mathsf{AC}^{k}\subseteq\textsf{TC}^{k}\subseteq\textsf{NC}^{k+1}.$
@@
$\textsf{NC}^{0}\subsetneq\mathsf{AC}^{0}\subsetneq\textsf{TC}^{0}\subseteq\textsf{NC}^{1}\subseteq\mathsf{L}\subseteq\textsf{NL}\subseteq\textsf{SAC}^{1}\subseteq\mathsf{AC}^{1}.$
We note that functions that are $\textsf{NC}^{0}$-computable can only depend on a bounded number of input bits. Thus, $\textsf{NC}^{0}$ is unable to compute the AND function. It is a classical result that $\mathsf{AC}^{0}$ is unable to compute Parity [Smo87]. The containment $\textsf{TC}^{0}\subseteq\textsf{NC}^{1}$ (and hence, $\textsf{TC}^{k}\subseteq\textsf{NC}^{k+1}$) follows from the fact that $\textsf{NC}^{1}$ can simulate the Majority gate.
@@
Further circuit classes. The complexity class $\mathsf{MAC^{0}}$ is the set of languages decidable by constant-depth uniform circuit families with a polynomial number of AND, OR, and NOT gates, and a single Majority gate at the output. The class $\mathsf{MAC^{0}}$ was introduced (but not so named) in [ABFR91], where it was shown that $\mathsf{MAC^{0}}\subsetneq\textsf{TC}^{0}$. This class was subsequently given the name $\mathsf{MAC^{0}}$ in [JKS02].
@@
The complexity class FOLL is the set of languages decidable by uniform $\mathsf{AC}$ circuit families of depth $O(\log\log n)$ and polynomial size. It is known that $\mathsf{AC}^{0}\subsetneq\textsf{FOLL}\subsetneq\mathsf{AC}^{1}$, and it is open as to whether FOLL is contained in NL [BKLM01].
@@
We will be particularly interested in NC circuits of quasipolynomial size (i. e., $2^{O(\log^{k}n)}$ for some constant $k$). For a circuit class $\mathcal{C}\subseteq\textsf{NC}$, the analogous class permitting a quasipolynomial number of gates is denoted $\textsf{quasi}\mathcal{C}$. We will focus specifically on $\mathsf{quasiAC^{0}}$. Here, we stress that our results for $\mathsf{quasiAC^{0}}$ will be stated for the nonuniform setting. Note that $\mathsf{DLOGTIME}$ uniformity does not make sense for $\mathsf{quasiAC^{0}}$, as we cannot encode gate indices using $O(\log n)$ bits. Nonetheless, there exist suitable notions of uniformity for $\mathsf{quasiAC^{0}}$ [Bar92, FGSTT20].
@@
Bounded nondeterminism. For a complexity class $\mathcal{C}$, we define $\exists^{\log^{i}n}\mathcal{C}$ to be the set of languages $L$ such that there exists an $L^{\prime}\in\mathcal{C}$ such that $x\in L$ if and only if there exists $y$ of length at most $O(\log^{i}|x|)$ such that $(x,y)\in L^{\prime}$. Similarly, define $\forall^{\log^{i}n}\mathcal{C}$ to be the set of languages $L$ such that there exists an $L^{\prime}\in\mathcal{C}$ such that $x\in L$ if and only if for all $y$ of length at most $O(\log^{i}|x|)$, $(x,y)\in L^{\prime}$. For any $i\geq 0$ and any $c\geq 0$, both $\exists^{\log^{i}n}\mathsf{FOLL}$ and $\forall^{\log^{i}n}\mathsf{FOLL}$ are contained in $\textsf{quasi}\mathsf{FOLL}$, and so cannot compute Parity [CTW13, Smo87]. Note that $\forall^{\log n}\mathcal{C}\cup\exists^{\log n}\mathcal{C}\subseteq\mathsf{AC^{0}}(\mathcal{C})$.
@@
Time and space-restricted Turing machines. When considering complexity classes defined by Turing machines with a time bound $t(n)\in o(n)$, we use Turing machines with random access and a separate address (or index) tape. After writing an address, the machine can go to a query state reading the symbol from the input at the location specified by the address tape. As usual, the machines are otherwise allowed to have multiple work tapes.
@@
For functions $t(n),s(n)\in\Omega(\log n)$, the classes $\mathsf{DTISP}(t(n),s(n))$ and $\mathsf{NTISP}(t(n),s(n))$ are defined to consist of decision problems computable by deterministic (resp. nondeterministic) $t(n)$ time and $s(n)$ space bounded Turing machines. Be aware that there must be one Turing machine that simultaneously satisfies the time and space bound. For details we refer to [Vol99, Section 2.6]. For further reading on the connection to $\mathsf{quasiAC^{0}}$, we refer to [Bar92, FGSTT20].
@@
We frequently use $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ and $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. However, because of how small the time and space bounds are for these classes, when we abuse notation to say some function (not necessarily decision problem) is computable in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, there is some ambiguity as to what this might mean. We use the following two equivalent definitions.
[…]
@@
[…]
@@
Proof.
($\Rightarrow$) Suppose $f$ is computable by a Turing machine $M$ using $\operatorname{poly}(\log n)$ time and $O(\log n)$ workspace (including the output). Then the following machine $N$ computes the function $(x,i)\mapsto f(x)_{i}$. On input $x,i$, $N$ simulates $M(x)$ while leaving $i$ untouched. After the simulation has completed, one of $N$’s worktapes contains $f(x)_{i}$, while $i$ is still on the input tape. Next, using another worktape to keep a counter, $N$ walks through $f(x)$ until it gets to the $i$-th bit, and outputs that bit of $f(x)$.
The entirety of the computation after the simulation of $N$ uses an amount of space that is $|i|\leq O(|f(x)|)\leq O(\log|x|)$, and time that is at most $|i|^{2}\leq O(\log^{2}|x|)$ (in fact, by noting that incrementing a counter by $+1$ $k$ many times only incurs $O(k)$ many changes to the bits of the counter, this can be reduced to $O(\log|x|)$ as well).
($\Leftarrow$) Suppose the function $(x,i)\mapsto f(x)_{i}$ is computable by a Turing machine $N$ using $\operatorname{poly}(\log n)=\operatorname{poly}(\log|x|)$ time and $O(\log n)=O(\log|x|)$ space. The following machine $M$ computes $f$ in the manner in which, when started with $x$ on its input tape, at the end of the computation $f(x)$ is the only thing written on the work tape. On input $x$, $M$ simulates $N(x,0),N(x,1),\dotsc$, writing those bits in the specified place on its work tape, until it reaches an $i$ such that $N(x,i)=\bot$. As $N$ was called $|f(x)|+1$ times, this multiplies the time complexity of $N$ by $|f(x)|\leq O(\log|x|)$. However, the space need does not increase except to keep track of the integer $i$ for which the machine will next simulate $N(x,i)$, plus the space for the output, as the space to simulate $N(x,i)$ can be re-used when simulating $N(x,i+1)$. Thus the space used is that of $N$, plus the space for the output, plus the space to maintain $i$, which is $|i|\leq\log(|f(x)|+1)\leq O(\log\log|x|)$, so the total space is still $O(\log|x|+\log\log|x|)=O(\log|x|)$. The additional clearing and marking at the end at most doubles the time and adds a constant amount of space usage. ∎
Fact 2.4.
@@
If there are no non-deterministic guesses made, i.e. our machine $M$ was deterministic, then we can get a CNF from the above decision tree by using De Morgan’s law: negate the leaves of the decision tree, get the corresponding DNF of the negated tree, then negate the DNF. (This does not work in the nondeterminstic case, as the disjunction includes all inputs such that there exists nondeterministic guesses that lead to an accepting leaf, and when we negate the leaves we end up with co-nondeterminism instead of non-determinism.) ∎
@@
Proof.
@@
$\mathop{\mbox{{Reach}}}\nolimits(\alpha,\beta,0)$ $\iff\beta\text{ is reachable from $\alpha$ in at most one computation step}$
@@
[…]
@@
[…]
@@
Disjunctive truth-table reductions. We finally recall the notion of a disjunctive truth-table reduction. Again let $L_{1},L_{2}$ be languages. We say that $L_{1}$ is disjunctive truth-table (dtt) reducible to $L_{2}$, denoted $L_{1}\leq_{dtt}L_{2}$, if there exists a function $g$ mapping a string $x$ to a tuple of strings $(y_{1},\ldots,y_{k})$ such that $x\in L_{1}$ if and only if there is some $i\in\{1,\dots,k\}$ such that $y_{i}\in L_{2}$. When $g$ is computable in a complexity class $\mathcal{C}$, we call this a $\mathcal{C}$-dtt reduction, and write $L_{1}\leq_{dtt}^{\mathcal{C}}L_{2}$. For any class $\mathcal{C}$, $\mathcal{C}$-dtt reductions are intermediate in strength between $\mathcal{C}$-many-one reductions and $\mathcal{C}$-Turing reductions.
Fact 2.7.
@@
Proof.
Let $L\in\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$, and suppose $L^{\prime}\leq_{dtt}^{\mathsf{AC^{0}}}L$ via $g$. We show that $L^{\prime}$ is also in $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$. Let $V\in\mathsf{AC^{0}}(\mathcal{C})$ be a predicate such that $x\in L\iff[\exists^{\log^{2}|x|}w]\,V(x,w)$ for all strings $x$. Then we have $x\in L^{\prime}$ iff $\exists\,i\in\{1,\dots,k\},\exists\,w\in\{0,1\}^{\log^{2}n}:V(g(x)_{i},w)$, where $g(x)_{i}$ denotes the $i$-th string output by $g(x)=(y_{1},\dotsc,y_{k})$. Since $g$ is, in particular, polynomial size, we have $k\leq\operatorname{poly}(|x|)$, so we have the bit-length of the index $i$ is at most $O(\log|x|)$. Finally, since $\mathsf{AC^{0}}(\mathcal{C})$ denotes the oracle class, it is closed under $\mathsf{AC^{0}}$ reductions, and thus the function $(x,w,i)\mapsto V(g(x)_{i},w)$ is in $\mathsf{AC^{0}}(\mathcal{C})$, and therefore $L^{\prime}\in\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$. ∎

### v5

@@ 2.2 Computational Complexity
We assume that the reader is familiar with standard complexity classes such as $\mathsf{L}$, $\mathsf{NL}$, $\mathsf{NP}$, and $\mathsf{EXP}$. For a standard reference on circuit complexity, see [57]. We consider Boolean circuits using the $\mathsf{AND},\mathsf{OR},\mathsf{NOT},$ and $\mathsf{Majority},$ where $\mathsf{Majority}(x_{1},\ldots,x_{n})=1$ if and only if $\geq n/2$ of the inputs are $1$. Otherwise, $\mathsf{Majority}(x_{1},\ldots,x_{n})=0$. In this paper, we will consider $\mathsf{DLOGTIME}$-uniform circuit families $(C_{n})_{n\in\mathbb{N}}$. For this, one encodes the gates of each circuit $C_{n}$ by bit strings of length $O(\log n)$. Then the circuit family $(C_{n})_{n\geq 0}$ is called $\mathsf{DLOGTIME}$-uniform if (i) there exists a deterministic Turing machine that computes for a given gate $u\in\{0,1\}^{*}$ of $C_{n}$ ($|u|\in O(\log n)$) in time $O(\log n)$ the type of gate $u$, where the types are $x_{1},\ldots,x_{n}$, $\mathsf{NOT},\mathsf{AND},\mathsf{OR},\mathsf{Majority},$ or oracle gates, and (ii) there exists a deterministic Turing machine that decides for two given gates $u,v\in\{0,1\}^{*}$ of $C_{n}$ ($|u|,|v|\in O(\log n)$) and a binary encoded integer $i$ with $O(\log n)$ many bits in time $O(\log n)$ whether $u$ is the $i$-th input gate for $v$.
@@
Fix $k\geq 0$. We say that a language $L$ belongs to (uniform) $\mathsf{NC}^{k}$ if there exist a (uniform) family of circuits $(C_{n})_{n\in\mathbb{N}}$ over the $\mathsf{AND},\mathsf{OR},\mathsf{NOT}$ gates such that the following hold:
  • The $\mathsf{AND}$ and $\mathsf{OR}$ gates take exactly $2$ inputs. That is, they have fan-in $2$.
@@
The complexity class $\mathsf{AC}^{k}$ is defined analogously as $\mathsf{NC}^{k}$, except that the $\mathsf{AND},\mathsf{OR}$ gates are permitted to have unbounded fan-in. That is, a single $\mathsf{AND}$ gate can compute an arbitrary conjunction, and a single $\mathsf{OR}$ gate can compute an arbitrary disjunction. The class $\mathsf{SAC}^{k}$ is defined analogously, in which the $\mathsf{OR}$ gates have unbounded fan-in but the $\mathsf{AND}$ gates must have fan-in $2$. The complexity class $\mathsf{TC}^{k}$ is defined analogously as $\mathsf{AC}^{k}$, except that our circuits are now permitted $\mathsf{Majority}$ gates of unbounded fan-in. We also allow circuits to compute functions by using multiple output gates.
Furthermore, for a language $L$ the class $\mathsf{AC}^{k}(L)$, apart from Boolean gates, also allows oracle gates for $L$ (an oracle gate outputs $1$ if and only if its input is in $L$). If $K\in\mathsf{AC}^{k}(L)$, then $K$ is said to be $\mathsf{AC}^{k}$-Turing reducible to $L$. Finally, for some complexity class $\mathcal{C}$ denote $\mathsf{AC}^{k}(\mathcal{C})$ to be the set of decision problems that are $\mathsf{AC}^{k}$-Turing reducible to problems in $\mathcal{C}$. Be aware that here we follow the notation of [57], which is different from [58, 26] (where $\mathsf{AC}^{k}(\mathcal{C})$ is used to denote composition of functions).
@@
$\mathsf{NC}^{k}\subseteq\mathsf{SAC}^{k}\subseteq\mathsf{AC}^{k}\subseteq\mathsf{TC}^{k}\subseteq\mathsf{NC}^{k+1}.$
@@
$\mathsf{NC}^{0}\subsetneq\mathsf{AC}^{0}\subsetneq\mathsf{TC}^{0}\subseteq\mathsf{NC}^{1}\subseteq\mathsf{L}\subseteq\mathsf{NL}\subseteq\mathsf{SAC}^{1}\subseteq\mathsf{AC}^{1}.$
We note that functions that are $\mathsf{NC}^{0}$-computable can only depend on a bounded number of input bits. Thus, $\mathsf{NC}^{0}$ is unable to compute the $\mathsf{AND}$ function. It is a classical result that $\mathsf{AC}^{0}$ is unable to compute Parity [52]. The containment $\mathsf{TC}^{0}\subseteq\mathsf{NC}^{1}$ (and hence, $\mathsf{TC}^{k}\subseteq\mathsf{NC}^{k+1}$) follows from the fact that $\mathsf{NC}^{1}$ can simulate the $\mathsf{Majority}$ gate.
@@
We also use the following easy lemma:
Lemma 2.3.
The prime factors of a $O(\log n)$-bit integer can be computed in $\mathsf{AC^{0}}$.
(We remind the reader that we always refer to the uniform classes unless otherwise specified; without uniformity the result would be immediate as all functions of $\log n$ bits are in nonuniform $\mathsf{AC^{0}}$.)
Proof 2.4.
Each prime divisor of $n$ can be represented using $O(\log n)$ bits. As the numbers involved are only $O(\log n)$ bits, ordinary arithmetic function of these numbers can be computed in $\mathsf{AC}^{0}$, in particular, testing if an $O(\log n)$-bit number is prime, testing if one $O(\log n)$-bit number divides another. So, in parallel, for all numbers $x=2,3,\dotsc,n/2$, an $\mathsf{AC}^{0}$ circuit checks which ones are prime and divide $n$.
Further circuit classes.
@@
The complexity class $\mathsf{MAC^{0}}$ is the set of languages decidable by constant-depth uniform circuit families with a polynomial number of $\mathsf{AND},\mathsf{OR},$ and $\mathsf{NOT}$ gates, and a single $\mathsf{Majority}$ gate at the output. The class $\mathsf{MAC^{0}}$ was introduced (but not so named) in [4], where it was shown that $\mathsf{MAC^{0}}\subsetneq\mathsf{TC}^{0}$. This class was subsequently given the name $\mathsf{MAC^{0}}$ in [32].
@@
The complexity class $\mathsf{FOLL}$ is the set of languages decidable by uniform $\mathsf{AC}$ circuit families of depth $O(\log\log n)$ and polynomial size. It is known that $\mathsf{AC}^{0}\subsetneq\mathsf{FOLL}\subsetneq\mathsf{AC}^{1}$, and it is open as to whether $\mathsf{FOLL}$ is contained in $\mathsf{NL}$ [10].
@@
We will be particularly interested in $\mathsf{NC}$ circuits of quasipolynomial size (i. e., $2^{O(\log^{k}n)}$ for some constant $k$). For a circuit class $\mathcal{C}\subseteq\mathsf{NC}$, the analogous class permitting a quasipolynomial number of gates is denoted $\mathsf{quasi}\mathcal{C}$. We will focus specifically on $\mathsf{quasiAC^{0}}$. Here, we stress that our results for $\mathsf{quasiAC^{0}}$ will be stated for the nonuniform setting. Note that $\mathsf{DLOGTIME}$ uniformity does not make sense for $\mathsf{quasiAC^{0}}$, as we cannot encode gate indices using $O(\log n)$ bits. Nonetheless, there exist suitable notions of uniformity for $\mathsf{quasiAC^{0}}$ [9, 24].
Bounded nondeterminism.
@@
For a complexity class $\mathcal{C}$, we define $\exists^{\log^{i}n}\mathcal{C}$ to be the set of languages $L$ such that there exists an $L^{\prime}\in\mathcal{C}$ such that $x\in L$ if and only if there exists $y$ of length at most $O(\log^{i}|x|)$ such that $(x,y)\in L^{\prime}$. Similarly, define $\forall^{\log^{i}n}\mathcal{C}$ to be the set of languages $L$ such that there exists an $L^{\prime}\in\mathcal{C}$ such that $x\in L$ if and only if for all $y$ of length at most $O(\log^{i}|x|)$, $(x,y)\in L^{\prime}$. For any $i\geq 0$ and any $c\geq 0$, both $\exists^{\log^{i}n}\mathsf{FOLL}$ and $\forall^{\log^{i}n}\mathsf{FOLL}$ are contained in $\mathsf{quasi}\mathsf{FOLL}$, and so cannot compute Parity [17, 52]. Note that $\forall^{\log n}\mathcal{C}\cup\exists^{\log n}\mathcal{C}\subseteq\mathsf{AC^{0}}(\mathcal{C})$.
Time and space-restricted Turing machines.
@@
When considering complexity classes defined by Turing machines with a time bound $t(n)\in o(n)$, we use Turing machines with random access and a separate address (or index) tape. After writing an address, the machine can go to a query state reading the symbol from the input at the location specified by the address tape. As usual, the machines are otherwise allowed to have multiple work tapes.
@@
For functions $t(n),s(n)\in\Omega(\log n)$, the classes $\mathsf{DTISP}(t(n),s(n))$ and $\mathsf{NTISP}(t(n),s(n))$ are defined to consist of decision problems computable by deterministic (resp. nondeterministic) $t(n)$ time and $s(n)$ space bounded Turing machines. Be aware that there must be one Turing machine that simultaneously satisfies the time and space bound. For details we refer to [57, Section 2.6]. For further reading on the connection to $\mathsf{quasiAC^{0}}$, we refer to [9, 24].
@@
We frequently use $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ and $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. However, because of how small the time and space bounds are for these classes, when we abuse notation to say some function (not necessarily decision problem) is computable in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, there is some ambiguity as to what this might mean. We use the following two equivalent definitions. First we need some setup. For a function $f$ from bit-strings to bit-strings, we define its bit function $\text{bit-}f$ as follows. Write $f(x)_{i}$ to denote the $i$-th bit of $f(x)$. Then define
$\text{bit-}f(x,i)=\begin{cases}f(x)_{i}&i\leq|f(x)|\\ \bot&i>|f(x)|.\end{cases}$
Note that $\text{bit-}f$ is always a function which we may take to have at most two output bits (e.g., by encoding $f(x)_{i}$ by repeating the bit, and encoding $\bot$ by $01$), i.e. a pair of decision problems.
@@
    The two bits of $\text{bit-}f$ are each decision problems in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, i. e., can each be decided by a Turing machine using $\operatorname{poly}(\log n)$ time and $O(\log n)$ space.
@@
Proof 2.6.
($\Rightarrow$) Suppose $f$ is computable by a Turing machine $M$ using $\operatorname{poly}(\log n)$ time and $O(\log n)$ workspace (including the output). Then the following machine $N$ computes the function $(x,i)\mapsto f(x)_{i}$. On input $(x,i)$, NN simulates $M(x)$ while leaving $i$ untouched. After the simulation has completed, one of $N$’s worktapes contains $f(x)_{i}$, while $i$ is still on the input tape. Next, using another worktape to keep a counter, $N$ walks through $f(x)$ until it gets to the $i$-th bit, and outputs that bit of $f(x)$.
The entirety of the computation after the simulation of $N$ uses an amount of space that is $|i|\leq O(|f(x)|)\leq O(\log|x|)$, and time that is at most $|i|^{2}\leq O(\log^{2}|x|)$ (in fact, by noting that incrementing a counter by +1+1 $k$ many times only incurs $O(k)$ many changes to the bits of the counter, this can be reduced to $O(\log|x|)$ as well).
($\Leftarrow$) Suppose the function $(x,i)\mapsto f(x)_{i}$ is computable by a Turing machine $N$ using $\operatorname{poly}(\log n)=\operatorname{poly}(\log|x|)$ time and $O(\log n)=O(\log|x|)$ space. The following machine $M$ computes $f$ in the manner in which, when started with $x$ on its input tape, at the end of the computation $f(x)$ is the only thing written on the work tape. On input $x$, $M$ simulates $N(x,0),N(x,1),\dotsc$, writing those bits in the specified place on its work tape, until it reaches an $i$ such that $N(x,i)=\bot$. As $N$ was called $|f(x)|+1$ times, this multiplies the time complexity of $N$ by $|f(x)|\leq O(\log|x|)$. However, the space need does not increase except to keep track of the integer $i$ for which the machine will next simulate $N(x,i)$, plus the space for the output, as the space to simulate $N(x,i)$ can be re-used when simulating $N(x,i+1)$. Thus the space used is that of $N$, plus the space for the output, plus the space to maintain $i$, which is $|i|\leq\log(|f(x)|+1)\leq O(\log\log|x|)$, so the total space is still $O(\log|x|+\log\log|x|)=O(\log|x|)$. The additional clearing and marking at the end at most doubles the time and adds a constant amount of space usage.
Fact 1.
@@
If there are no non-deterministic guesses made, i.e. our machine $M$ was deterministic, then we can get a CNF from the above decision tree by using De Morgan’s law: negate the leaves of the decision tree, get the corresponding DNF of the negated tree, then negate the DNF. (This does not work in the nondeterminstic case, as the disjunction includes all inputs such that there exists nondeterministic guesses that lead to an accepting leaf, and when we negate the leaves we end up with co-nondeterminism instead of non-determinism.)
@@
Proof 2.9.
This is essentially the proof of Savitch’s Theorem: A configuration of a Turing machine consist of the current state, the work and index tape content, but not the content of the input tape. For configurations $\alpha,\beta$ define the $\mathop{\mbox{{Reach}}}\nolimits$ predicate as follows:
@@
for $i\geq 1$. Thus, $\mathop{\mbox{{Reach}}}\nolimits(\alpha,\beta,i)$ holds if and only if $\beta$ can be reached from $\alpha$ in at most $2^{i}$ computation steps. Since the running time is bounded by $\log^{k}n$ for some $k$, by letting $\mathop{\mbox{{Start}}}\nolimits(w)$ denote the initial configuration for the input $w\in\Sigma^{*}$ and $\mathop{\mbox{{Accept}}}\nolimits$ the accepting configuration of $M$ (which can be assumed to be unique after a suitable manipulation of $M$), we have
@@
Now, it remains to observe that the inductive definition of the $\mathop{\mbox{{Reach}}}\nolimits$ predicate can be evaluated in $\mathsf{FOLL}$ since the recursion depth is $O(\log\log n)$, each recursion step is clearly in $\mathsf{AC^{0}}$ and there are only polynomially many configurations (because of the space bound of $\log n$).
@@
Disjunctive truth-table reductions.
We finally recall the notion of a disjunctive truth-table reduction. Again let $L_{1},L_{2}$ be languages. We say that $L_{1}$ is disjunctive truth-table (dtt) reducible to $L_{2}$, denoted $L_{1}\leq_{dtt}L_{2}$, if there exists a function $g$ mapping a string $x$ to a tuple of strings $(y_{1},\ldots,y_{k})$ such that $x\in L_{1}$ if and only if there is some $i\in\{1,\dots,k\}$ such that $y_{i}\in L_{2}$. When $g$ is computable in a complexity class $\mathcal{C}$, we call this a $\mathcal{C}$-dtt reduction, and write $L_{1}\leq_{dtt}^{\mathcal{C}}L_{2}$. For any class $\mathcal{C}$, $\mathcal{C}$-dtt reductions are intermediate in strength between $\mathcal{C}$-many-one reductions and $\mathcal{C}$-Turing reductions.
Fact 3.
@@
Proof 2.10.
Let $L\in\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$, and suppose $L^{\prime}\leq_{dtt}^{\mathsf{AC^{0}}}L$ via $g$. We show that $L^{\prime}$ is also in $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$. Let $V\in\mathsf{AC^{0}}(\mathcal{C})$ be a predicate such that $x\in L\iff[\exists^{\log^{2}|x|}w]\,V(x,w)$ for all strings $x$. Then we have $x\in L^{\prime}$ iff $\exists\,i\in\{1,\dots,k\},\exists\,w\in\{0,1\}^{\log^{2}n}:V(g(x)_{i},w)$, where $g(x)_{i}$ denotes the $i$-th string output by $g(x)=(y_{1},\dotsc,y_{k})$. Since $g$ is, in particular, polynomial size, we have $k\leq\operatorname{poly}(|x|)$, so we have the bit-length of the index $i$ is at most $O(\log|x|)$. Finally, since $\mathsf{AC^{0}}(\mathcal{C})$ denotes the oracle class, it is closed under $\mathsf{AC^{0}}$ reductions, and thus the function $(x,w,i)\mapsto V(g(x)_{i},w)$ is in $\mathsf{AC^{0}}(\mathcal{C})$, and therefore $L^{\prime}\in\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$.